A NEGATIVE ION PHOTOELECTRON SPECTROSCOPIC AND COMPUTATIONAL STUDY OF CrV AND MoV

\underline{BEAU J. BARKER}, SEAN M. CASEY, SIMSON ALEX, DOREEN G. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

The heteronuclear diatomic metal ions CrV^{−} and MoV^{−} were investigated using negative ion photoelectron spectroscopy. These bare, heteronuclear group 5 and 6 transition metal dimers allow the study of multiple metal-metal bonding free of ligand effects. The photoelectron spectra, obtained at 488 nm with an instrumental resolution of about 5 meV (40 cm^{−1}), provide measurements of the electron affinities, vibrational frequencies for both the anion and neutral states, bond length changes upon electron detachment, excited electronic state energies and spin-orbit splittings. The CrV^{−} and MoV^{−} spectra display transitions to the multiply-bonded ^{2}Δ (dπ)^{4}(dδ)^{3}(dσ)^{2}(sσ)^{2} ground states of the neutral molecules and to several excited states. Addition of an electron to the vacant σ^{∗} orbital gives the ^{3}Δ anion, which is found to be the ground state of CrV^{−} but a low lying excited state of MoV^{−}. Addition of a dδ electron yields the ^{1}Σ^{+} MoV^{−} ground state. Density functional theory calculations of the anions and neutrals were performed to help elucidate the spectroscopic assignments and observed periodic trends.